OpenStudy (anonymous):

Determine the direction in which the parabola opens, and the vertex. y2 = + 6x + 14 A. Opens upward; ( -3, 5) B. Opens upward; ( 3, 5) C. Opens to the right; ( 5, 3) D. Opens to the right; ( 5, -3) its NOT D because I put D and got it wrong..

OpenStudy (lgbasallote):

uhmm is that supposed to be \[\large y^2 = x^2 + 6x + 14\] or something?

OpenStudy (anonymous):

y2 = + 6x + 14 is what the question says.. im not sure..

OpenStudy (phi):

looks like a typo in the question

OpenStudy (mathmate):

C would correspond to x=y^2-6y+14 A and B cannot be right if there is a y^2 term.

OpenStudy (anonymous):

ok. thnx.